Session: Feb 2002 (session 1)
Wavelength: 18. cm


To give feedback on this experiment, click here

Experiment Feedback

Torun (Kazimierz Borkowski at Wed Feb 20 11:00:31 MET 2002) Success
ES042C. No known problems (remarks for the wavelength do not apply).

Effelsberg (A. Kraus at Wed Feb 20 13:30:18 MET 2002) Minor failures
Data lost between UT0312-0329: wrong cable wrap accidentally chosen by the operator.

Jodrell_Bank (Alastair Gunn at Fri Feb 22 10:09:49 MET 2002) Minor failures
ES042C. Telescope parked due to high winds between 00:24 and 01:36 UT and again from 07:36 UT until end of experiment. No other known problems.

Cambridge (Alastair Gunn at Fri Feb 22 10:10:16 MET 2002) Failure
ES042C. Cambridge did not observe due to serious fault with recorder.

Westerbork (Tony Foley at Fri Feb 22 16:07:00 MET 2002) Minor failures
no known problems

Onsala (John Conway at Tue Feb 26 17:07:15 MET 2002) Minor failures
The EVN spare capstan motor arrived and was installed. Did record tests but no time for a new head-cal, subsequent head cal the next day showed forward reverse offset close to that in head.ctl file, perhap 10 microns out. Did not observe last hour of expt in order to perform a head cal before next expt.

Medicina (G.Maccaferri at Wed Mar 6 10:28:26 MET 2002) Success